### v2

# Finding the Fairest Voting System Using Likelihood Analysis

Abstract
This paper takes a statistical approach to determine which of 10 voting systems, or preference aggregation rules, is the fairest based on their probabilistic likelihood of violating Arrow’s five social choice criteria. The voting systems considered are: Plurality, Borda, Dowdall, Top Two (Plurality Runoff), Instant Runoff, Coombs, Baldwin, Copeland, Pairwise Majority, and Minimax. This paper builds upon the work of Dougherty and Heckelman (2020) by computing violation frequencies for elections with a greater number of alternatives. Elections with up to 50,000 voters and between three and six alternatives are simulated using both Impartial Culture and Impartial Anonymous Culture. The results of these simulations produce new IIA violation likelihoods for each method and show that Pairwise Majority is the most likely to jointly satisfy all five of Arrow’s criteria. Furthermore, of the systems that satisfy transitivity, the Baldwin method is most likely to jointly satisfy all five of Arrow’s criteria in elections with three alternatives. As the number of alternatives increase, the joint satisfaction frequencies decrease rapidly for all systems.
1 Introduction
Voting systems can be evaluated based on whether they satisfy or violate social choice criteria. This paper takes a statistical approach to determine which voting system is the fairest based on the probabilistic likelihood of violating Arrow’s five social choice criteria.
In total, 10 voting systems are analyzed with respect to five social choice criteria. Specifically, the methods considered are Plurality, Borda, Dowdall, Top Two (Plurality Runoff), Instant Runoff, Coombs, Baldwin, Copeland, Pairwise Majority, and Minimax. Through simulations with up to 50,000 voters and six alternatives, the violation frequencies are computed for each method. To model voters’ preferences, two cultures are used: Impartial Culture (IC) and Impartial Anonymous Culture (IAC). First, voting systems are tested with respect to each individual criteria. Then, for each voting system, Joint-Satisfaction (JS) frequencies, the probability that systems jointly satisfy all five conditions, are computed. These frequencies are used to determine the fairest system.
This paper builds upon the work of previous empirical studies in social choice theory. Compared to prior works, this study computes violation frequencies for elections with a greater number of voters and alternatives than most literature. For the case of three-alternatives elections, the majority of our results align with Dougherty and Heckelman (2020). This paper both extends their analysis for up to six alternatives and finds higher lower bounds for the frequency of IIA violations. The main findings are summarized below:
Result 1: Of the 10 voting systems analyzed, Pairwise Majority is the most likely to jointly satisfy Arrow’s five social choice criteria. In elections with three alternatives and at least 100 voters, Pairwise Majority has Joint-Satisfaction frequencies of over 85%.
Result 2: Of the systems that satisfy transitivity, Baldwin’s Method has the highest Joint-Satisfaction frequencies. The frequencies, however, are less than 20%.
Result 3: As the number of alternatives increases, the Joint-Satisfaction frequencies decrease for each voting method. For all systems except Pairwise Majority, the Joint-Satisfaction frequencies in elections with at least 30 voters and four alternatives are 0%.
The paper proceeds as follows: Sections 2 and 3 detail all required background knowledge. Section 4 contains this study’s contributions and new results. Section 5 describes the experiment set-up and methodology. Section 6 presents the results, and Section 7 discusses the differences in results to past papers. Section 8 concludes the paper and highlights the key points.
2 Voting Systems
We follow the same assumptions as in Dougherty and Heckelman (2020). We define below the voting systems analyzed in this paper which are a subset of those considered in their paper. In Dougherty and Heckelman (2020), Plurality, Minimax, Top Two, and Instant Runoff are said to be part of Group 1 due to their similar numerical results, and we adopt this convention here.
[…]
Plurality -
Alternatives are ranked based on their first-place vote tally.
Borda Count -
Let the election have $n$ alternatives. For each 1st place vote, alternatives receive $n$ points, for each 2nd place vote they receive $n-1$ points, and so on until for an $n$th place vote, alternatives receive one point. To obtain a societal preference order, the alternatives are ranked based on their point total.
Dowdall -
Let the election have $n$ alternatives. For each first place vote, alternatives receive 1 point, for a second place vote they receive $\frac{1}{2}$ point, for a third place vote they receive $\frac{1}{3}$ point and so on until for an $n$th place vote, alternatives receive $\frac{1}{n}$ points. The voters are then ranked based on their point tally.
Top Two/Plurality Runoff -
There are two rounds. The first round is a plurality election. If a alternative receives a majority, they win the election. If not, the voters vote again, but can only chose between the two alternatives who receive the most first-place votes in the first round. The alternative with the most first-place votes in the second round wins the election.111 If there is a majority winner in the first round, alternatives are ranked based on their first-place vote tally. If there is no majority winner, the top two alternatives are compared and ranked first and second (or tied first) depending on their relative vote totals in the second round. Then the other alternatives are ranked below them based on their first round vote totals.
Instant Runoff/Hare -
In each round, the alternative(s) with the least first place votes is eliminated. The votes of the eliminated alternative(s) are transferred to the alternatives ranked below them. The process repeats until all alternatives are eliminated. Alternatives are ranked based on reverse order of the round they were eliminated.
Coombs -
[…]
Baldwin -
This system follows the same procedure as Instant Runoff, but has a different elimination rule. Each round, the alternative with the least points, as calculated by the Borda Count rule, is eliminated.
Pairwise Majority -
All alternatives are compared head to head pairwise. For two alternatives $A$ and $B$, if $A$ wins the head to head then $A\succ B$. If $B$ wins the head to head then $B\succ A$. If they tie then $A\sim B$. Based on the comparisons, a transitive ordering is formed if one exists.
Copeland/Pairwise Comparison -
All alternatives are compared head to head pairwise. For each comparison, the alternative that wins head to head receives 1 point. If the two alternatives tie they receive $\frac{1}{2}$ point each. After all comparisons are performed, alternatives are ranked based on their point total.
Minimax/Simpson-Kramer -
All alternatives are compared head to head. For each alternative, their score is the magnitude of their worst pairwise defeat. Alternatives are ranked in the reverse order of their score (with lower scores giving higher rank).
2.1 Arrow’s Criteria of Fairness
[…]
  1. 1.
    Non-Dictatorship (ND) - There is no individual whose preferences uniquely determine the social raking of all alternatives, regardless of how other individuals rank the alternatives.
  2. 2.
    Unrestricted Domain (UD) - The domain of the voting system can take any preference schedule as input. This means that for every preference schedule, the voting system produces societal preferences between any two alternatives; there are no restrictions on how voters can vote. Note that, for the scope of this paper, this condition does not imply the rankings have to be transitive.
  3. 3.
    Transitivity (TR) - The voting method always produces transitive preferences, i.e., the method always yields a societal preference order. Formally, for any three candidates $A$, $B$, and $C$, $A\succeq B$ and $B\succeq C$ implies $A\succeq C$.
  4. 4.
    (Pareto) Unanimity (U) - If all voters prefer alternative $A$ over alternative $B$, then in the societal preference order, $A\succ B$.
  5. 5.
[…]
3 Arrow’s Theorem
[…]
Table 1: Criteria satisfied by 10 Voting Systems
Voting System ND UD TR U IIA
Plurality
Borda Count
Instant Runoff
Coombs Method
Copeland
Top Two
Dowdall
Pairwise Majority
Minimax
Baldwin
[…]
4 New and Expanded Results
The numerical probability of violating each of Arrow’s five conditions is investigated in Dougherty and Heckelman (2020). The study also calculates the likelihood that voting systems violate at least one of Arrow’s conditions. The new contributions of this paper that extend their prior results are:
  1. 1.
    This paper collects more empirical data, analyzing elections for up to 50,000 voters and six alternatives.
  2. 2.
    This paper confirms and collates the results of previous studies, providing extensive data of the frequencies at which voting systems violate criteria.
  3. 3.
    This paper develops a new methodology to test for IIA violations, which finds substantially more violations than previous algorithms, and thus creates a new greatest lower bound for IIA violations.
  4. 4.
    A new metric, Joint-Satisfaction frequencies, is introduced and formalized to test for the fairness of voting systems. This is similar to Dougherty and Heckelman (2020) regarding the likelihood of violating at least one criteria.
The data tables with the results showing percentages of violations for each criterion and Joint-Satisfaction frequencies are located in Appendix A.
5 Simulations
All computational models were implemented using C++ and Python. The voting systems (Plurality, Borda, Instant Runoff, Coombs, Copeland, Dowdall, Pairwise Majority, Minimax, Top Two, and Baldwin) were tested against Arrow’s criterion (ND, UD, IIA, U, TR) for three to six alternatives.
Ties were handled as follows: For Top Two, there is a possibility of a tie in the first round, where more than two alternatives have equal number of votes. In this case, ties are broken randomly. In all other methods, if alternatives are tied, they have the same rank in the societal preference order.
[…]
For each setting (fixing the criterion, number of voters, number of alternatives, and culture) 10,000 preference schedules were generated. Then the violation frequency, the percentage of times a voting system violates the criterion, was calculated for each of the 10 voting systems. For JS Frequencies, this involved testing voting systems against multiple criteria consecutively. Due to the simulation process, the frequency of violations calculated are numerical probabilities, not theoretical.
All computations were carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC1920103.
6 Results and Analysis
The following sections details the likelihood of voting systems violating Arrow’s social choice criterion. Note that all voting methods considered satisfy Non-Dictatorship and Unrestricted Domain by definition and as such are excluded from the data.
6.1 Independence of Irrelevant Alternatives (IIA)
Key Observations in IIA Data
  1. 1.
    In elections with at least three alternatives and two voters, Group 1 systems (Plurality, Instant Runoff, Minimax, and Top Two) violate IIA more than 99.99% of the time.
  2. 2.
    As the number of voters increases, IIA violations generally increase under IC. For all systems that violate IIA, the frequency of violations approaches 100% as the number of voters increase. For elections with at least 100 voters and three alternatives, IIA is almost never satisfied. However, this trend is not always observed as many voting systems violate IIA at a lower frequency in elections with four voters than elections with three voters. The IAC data does not strictly abide by this trend.
  3. 3.
    As the number of alternatives increases, IIA violations increase. In elections with 4 alternatives and at least 100 voters, all systems violate IIA more than 99.7% of the time for both IC and IAC. For elections with at least five alternatives, IIA is almost never satisfied.
Tables 2 and 3 show percentage violations for three-alternative elections under IC and IAC. Figures 1 and 2 plot the IC and IAC data respectively. Note that Pairwise Majority cannot violate IIA, thus the violation frequencies are not computed.
Table 2: Percentage IIA Violations for 3 alternatives (IC Data)
Voting System Number of Voters
3 4 5 10 30 50 100 1000
Plurality 100 100 100 100 100 100 100 100
Instant Runoff 100 100 100 100 100 100 100 100
Top Two 100 100 100 100 100 100 100 100
Minimax 100 100 100 100 100 100 100 100
Copeland 97.35 73.63 99.95 99.99 100 100 100 100
Dowdall 97.35 99.52 92.87 99.79 100 100 100 100
Borda Count 97.35 73.63 92.87 94.79 99.97 99.99 100 100
Coombs 97.35 73.63 92.87 94.79 99.99 100 100 100
Baldwin 97.35 73.63 92.87 94.79 99.96 100 100 100
Figure 1: IIA Violations in elections with three alternatives (IC)
Under IC, in elections with at least 30 voters, the frequency of violating IIA is 100%. In elections with three voters, Dowdall, Copeland, Coombs, Borda, and Baldwin show identical IIA violations.
Table 3: Percentage IIA Violations for 3 alternatives (IAC Data)
Voting System Number of Voters
3 4 5 10 30 50 100 1000
Plurality 100 100 100 100 100 100 100 100
Instant Runoff 100 100 100 100 100 100 100 100
Top Two 100 100 100 100 100 100 100 100
Minimax 100 100 100 100 100 100 100 100
Copeland 89.08 56.66 97.3 97.75 99.8 99.99 100 99.99
Dowdall 89.08 95.03 78.62 91.8 96.72 97.29 96.91 97.03
Borda Count 89.08 56.66 78.62 74.29 84.52 84.54 84.14 83.64
Coombs 89.08 56.66 78.62 74.29 87.4 86.41 85.7 85.52
Baldwin 89.08 56.66 78.62 74.29 83.23 83.38 83.35 82.49
The IAC data has slightly more promising results and there is significant variation between systems. Coombs, Borda, and Baldwin all violate IIA less than 86% of the time, even for a sizable electorate. IIA violations appear to slightly decrease for Baldwin, Borda, and Coombs as the number of voters increases. The percentage violations for 1000 voters is less than that for 100 voters.
Figure 2: IIA Violations in elections with three alternatives (IAC)
Our three-alternative data substantially differs from Dougherty and Heckelman (2020). For Dowdall, Coombs, Copeland, and Baldwin our algorithm reports greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points. These differences are further explained in Section 7.
6.2 Pareto Unanimity (U)
In the following, a Pareto pair is defined to be a pair of alternatives where all voters prefer one alternative in the pair over the other.
Only Group 1 systems can theoretically violate U. Even for these systems, U is nearly always satisfied. In elections with at least 30 voters, all systems have a 0% probability of violating U. U is almost always satisfied because encountering a Pareto pair is very rare, especially as the number of voters increases. Table 4 shows the frequency with which Instant Runoff violates U.
Table 4: Percentage Unanimity Violations Instant Runoff
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 16.23 58.00 86.12 96.61
3 2.56 17.38 41.34 62.30
4 0.43 4.59 13.93 28.63
5 0.05 1.14 4.58 10.87
10 0.00 0.00 0.00 0.07
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 59.94 86.13 96.56
3 11.02 21.64 41.69 63.20
4 4.42 7.96 15.58 28.35
5 2.32 2.83 5.00 10.14
10 0.19 0.06 0.04 0.05
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
[…]
Table 5 shows U violations for elections with four alternatives using IAC for Group 1 systems. Figure 3 is a graph of the data.
Table 5: Percentage U Violations 4 alternatives IAC
[…]
1 2 3 4 5 10 30 50 100
Plurality/InstantRunoff/TopTwo 100.00 59.94 21.64 7.96 2.83 0.06 0 0 0
Minimax 100.00 44.53 12.93 4.04 0.98 0.01 0 0 0
Figure 3: U Violations in elections with four alternatives (IAC)
While unanimity does not pose issues in large elections, for election with few voters and many candidates, this criterion is violated often.
6.3 Transitivity (TR)
Pairwise Majority is the only system studied that violates transitivity. Table 6 shows the percentage TR violations for Pairwise Majority. Figure 4 is a 3D plot of the IC data.
Key Observations
  1. 1.
    As the number of alternatives increases, the transitivity violations increase.
  2. 2.
[…]
Table 6: Percentage Transitivity Violations Pairwise Majority
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
2 33.47 66.49 86.56 95.54
3 5.54 17.12 32.60 48.93
4 30.69 63.91 85.84 95.51
5 6.87 21.04 39.59 57.39
10 23.81 54.73 78.98 92.50
30 17.31 43.78 68.77 86.03
50 15.36 40.17 64.69 82.66
100 13.08 35.99 60.08 78.92
1000 9.95 29.11 51.63 71.18
5000 9.35 27.45 49.01 69.16
10000 9.19 27.36 48.90 68.60
30000 9.11 26.48 47.95 68.11
50000 8.93 26.23 47.96 67.61
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
2 28.81 63.97 86.10 95.35
3 3.57 14.97 31.66 48.87
4 23.65 60.59 85.05 95.49
5 4.77 19.17 38.71 57.56
10 15.50 48.88 77.53 92.31
30 9.58 35.85 66.47 85.58
50 8.19 32.14 61.44 82.22
100 7.23 28.51 56.13 78.15
1000 6.46 24.54 47.83 69.68
5000 6.24 24.03 46.76 67.73
10000 6.34 24.13 46.72 67.08
30000 6.45 24.06 46.92 66.98
50000 6.34 23.90 46.49 67.21
Figure 4: TR Violations in Pairwise Majority elections (IC)
[…]
6.4 Joint Satisfaction of the Conditions
Simulations were run to find the percentage of times a voting system violates at least one criteria. The order in which criteria were tested was U, IIA, and TR. The other criterion (ND, UD) were not simulated as no system violates ND or UD. Joint-Satisfaction (JS) Frequencies, defined as the percentage of times systems satisfied all criteria simultaneously, serve as an indication of fairness of a voting system.
For every system that violates IIA, the JS Frequencies almost exactly align with the likelihood of satisfying IIA. In elections with more than three alternatives and two voters, no system had a JS Frequency of 100%, which is consistent with Arrow’s Impossibility Theorem. Table 7 shows the results for Baldwin. All other data tables can be found in Appendix A.
Table 7: JS Frequencies (%) for Baldwin
(a) IC Data
Number of voters Number of alternatives
3       4       5 6
10 5.16 0.01 0.00 0.00
30 0.10 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3       4       5 6
10 25.16 0.25 0.01 0.00
30 16.87 0.27 0.00 0.00
50 16.79 0.28 0.00 0.00
100 17.14 0.16 0.00 0.00
1000 17.27 0.22 0.00 0.00
5000 16.63 0.27 0.00 0.00
10000 16.68 0.29 0.00 0.00
30000 17.07 0.22 0.00 0.00
50000 16.88 0.33 0.00 0.00
Table 8 shows a comparison of all the voting systems for 3 alternative elections under IAC and Figure 5 is a graph of the data.
Table 8: Joint-Satisfaction Frequencies for 3 alternatives (IAC)
Voting System Number of Voters
10 100 1000 5000 10000 50000
Group 1 0.00 0.00 0.00 0.00 0.00 0.00
Copeland 1.90 0.02 0.00 0.02 0.03 0.02
Dowdall 7.41 3.16 3.13 3.11 2.97 3.10
Borda Count 25.16 16.15 16.00 15.48 15.37 15.75
Coombs 25.16 14.50 14.20 13.90 13.40 14.02
Baldwin 25.16 17.14 17.27 16.63 16.68 16.88
Pairwise Majority 84.50 92.77 93.54 93.76 93.67 93.66
Figure 5: JS Frequencies in elections with three alternatives (IAC)
Figure 6: JS Frequencies for Black, Borda, Baldwin, Coombs
The data shows that Pairwise Majority is the most likely to satisfy all the criteria jointly for elections with three alternatives. As the number of voters increase in Pairwise Majority, the JS Frequencies increase, while for other systems the JS Frequencies either decrease or remain approximately constant. This is due to the fact that as the number of voters increase, transitivity violations become more rare, while, in general, IIA violations become more common.
There are also clear differences in performance in the other systems in three-alternative elections. Of the systems that satisfy transitivity, Borda Count, Baldwin, and Coombs have the highest JS Frequencies. Figure 6 shows a zoomed in version of Figure 5 for these three systems. For Borda, Baldwin, and Coombs the JS Frequencies also slightly increase as the number of voters becomes very large. This is because the IIA violations in these systems tend to marginally decrease as the number of voters increases from 30 to 50,000.
However, as the number of alternatives increases, the JS Frequencies decrease for every system. Figure 7 shows this trend clearly. In fact, for all systems excluding Pairwise Majority, the JS Frequencies for elections with at least 10 voters and 4 alternatives are less than 0.5%. This is because IIA is almost always violated in elections with at least four alternatives. Even for Pairwise Majority, the JS Frequencies sharply decrease as the number of alternatives increase, falling from 93.66% for three alternatives elections to 32.80% for six alternatives elections.
Figure 7: JS Frequencies in elections with 50,000 voters (IAC)
The conclusion for three-alternative elections is consistent with literature. In Dougherty and Heckelman (2020), the likelihood of violating at least one of Arrow’s conditions is computed. Their study agrees that Pairwise Majority has the greatest JS Frequencies. However, due to differences in the IIA data as discussed in Section 7, the data for Copeland, Coombs, Dowdall, Borda, and Baldwin does not match exactly. Most systems have lower JS Frequencies than in the 2020 study.
7 Remarks about Differences in IIA Results
The IIA algorithm used in this paper is different than that of Dougherty and Heckelman (2020) and we describe those differences here. We start by describing our algorithm.
As IIA is an inter-profile condition, to check if a voting system violates IIA, multiple preference schedules have to be compared. Our IIA algorithm loops through all possible pairs of alternative. For each alternative pair $A$ and $B$, the following test is done: First, find the societal preference between $A$ and $B$. Then, let $j$ equal the number of voters who prefer $A$ over $B$ and $k$ the number of voters who prefer $B$ over $A$. Then randomly generate 500 additional preference schedules in which $j$ voters still prefer $A$ over $B$ and $k$ voters prefer $B$ over $A$. The number 500 was chosen taking various factors into consideration including computational resources and program run time. If any of these new preference schedules produce a different societal preference between $A$ and $B$ than the original, an IIA violation occurs. This is repeated for every possible pair of alternatives.
This algorithm can miss violations, which is more likely when the number of voters is high. As the number of voters increases, there are more possible preference schedules where $j$ voters prefer $A$ over $B$ and $k$ voters vote $B$ over $A$. Thus, in the 500 new preference schedules generated, there may be no preference schedule that has a different societal order, even if one exists. Checking every single preference is impractical since there are $(\frac{c!}{2})^{v}$ possible schedules such that each voter maintains their relative preference between a pair, where $c$ is the number of alternatives and $v$ is the number of voters.
In Dougherty and Heckelman (2020), a different IIA algorithm was used. Our algorithm finds substantially more IIA violations that the algorithm in Dougherty and Heckelman (2020), in particular with the IAC data. For Dowdall, Coombs, Copeland, and Baldwin our algorithm reports a greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points.
7.0.1 Disagreement in IIA Conclusions under Copeland
Dougherty and Heckelman (2020) conclude that Copeland cannot violate IIA if a majority of voters have identical preferences. However, consider an election with three voters and three alternatives with the following preference schedule where a strict majority have preference order $A\succ B\succ C$:
2 0 0 1 0 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
Under Copeland, the societal preference order is $A\succ B\succ C$. Now, consider the pair of alternatives A and B. Suppose that one voter who has preference $A\succ_{i}B\succ_{i}C$ changes their preference to $C\succ_{i}A\succ_{i}B$. Now the preference schedule is:
1 0 0 1 1 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
The societal preference order is now $A=B=C$. Even though all voters kept their relative preference between A and B, the societal ranking between them changed. Copeland can violate IIA even if a majority have identical preferences.
Dougherty and Heckelman (2020) also state that Copeland violates IIA at a greater frequency for even number of voters $N$ than for the adjacent odd values $N-1$ and $N+1$. They claim that this is the case as there are more ties for elections with an even number of voters. While ties do make it easier to violate IIA, there are many cases for four voter elections where there are no ties and IIA is satisfied. Our analysis shows that for even number of voters, IIA violations in Copeland actually occur less frequently. For instance in three-alternative IC elections, Copeland violates IIA 97.35% of the time for three voters and only 73.63% of the time for four voters under IC.
Another, possibly related, disagreement was found while using their algorithm. For the purpose of comparing algorithms, we coded their IIA algorithm in our Python model for testing. Using their algorithm, we obtained the exact same data table that they did for both IC and IAC, with the exception of Copeland. For Copeland elections with three alternatives and at least four voters (in elections with three voters, the data is in agreement), their reported values are higher than what their algorithm produces. For instance, in elections with four voters and three alternatives under IC, simulations using their algorithm show a 73.7% chance of violating IIA, while the value reported in their paper is 88.9%. Notably, our IIA tables show 73.6% violations for identical elections.
This over-counting is also seen for IAC elections with more than four voters. For elections with 30 voters and three alternatives under IAC, our implementation of their algorithm yields 72% IIA violations, but the reported value is 78.5%. Note that as the number of voters increases, the extent of over-counting decreases. For IC, their data reaches 100% violations, and for IAC the difference between their values and our implementation of their algorithm becomes almost 0. The authors suspect an error in their Copeland method for greater than four voters.
8 Summary of Results
After comparing all systems, Pairwise Majority is the fairest system as it has the greatest likelihood of jointly satisfying all social choice criterion in elections with at least three alternatives and two voters. The other systems perform very poorly in comparison, primarily due to the extremely high frequency of IIA violations. Of the systems that satisfy transitivity, Baldwin, Borda, and Coombs perform the best, but have JS Frequencies of less than 17%.
Of all systems considered, Group 1 systems perform the worst. They violate IIA 100% of the time and as such, always have a 0% probability of jointly satisfying all criterion in elections with at least three alternatives.
Another key observation is that as the number of alternatives increases, the frequency with which voting systems violate criteria generally increases. For IIA, TR, and U, most systems increase in percentage violations as the number of alternatives increase. Thus, in elections with more alternatives, the JS Frequencies are significantly lower.
In the end, in alignment with Arrow’s Theorem, no voting system is perfect. Each method gives rise to results which are arguably undesirable from the point of view of a society. However, not all voting systems are equally flawed. While some systems always violate social choice criteria, there are systems with a high likelihood of satisfying all criteria. Finding and implementing such systems can help achieve fairer and more representative electoral outcomes.
9 Declarations
Competing Interests The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability Statement The raw data generated during this study is available upon request. All relevant processed data is included within this article and the appendix.
[…]
Appendix A Data Tables
A.1 IIA
Table 9: Percentage IIA Violations Plurality and Instant Runoff
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 99.99 100.00 100.00 100.00
Table 10: Percentage IIA Violations Top Two
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
Table 11: Percentage IIA Violations Minimax
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 99.99 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 99.99 100.00 100.00 100.00
Table 12: Percentage IIA Violations Borda
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 92.87 99.39 100.00 100.00
10 94.79 100.00 100.00 100.00
30 99.97 100.00 100.00 100.00
50 99.99 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.90 100.00 100.00
30 84.52 99.87 100.00 100.00
50 84.54 99.90 100.00 100.00
100 84.14 99.93 100.00 100.00
1000 83.64 99.91 100.00 100.00
5000 83.60 99.87 100.00 100.00
10000 84.12 99.90 100.00 100.00
30000 83.74 99.85 100.00 100.00
50000 83.96 99.76 100.00 100.00
Table 13: Percentage IIA Violations Coombs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 92.87 99.39 100.00 100.00
10 94.79 100.00 100.00 100.00
30 99.99 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.91 100.00 100.00
30 87.40 100.00 100.00 100.00
50 86.41 99.97 100.00 100.00
100 85.70 100.00 100.00 100.00
1000 85.52 99.99 100.00 100.00
5000 85.57 100.00 100.00 100.00
10000 85.91 99.99 100.00 100.00
30000 85.43 99.99 100.00 100.00
50000 85.45 99.98 100.00 100.00
Table 14: Percentage IIA Violations Dowdall
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 99.52 99.96 100.00 100.00
5 92.87 100.00 100.00 100.00
10 99.79 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 95.03 99.81 100.00 100.00
5 78.62 99.99 100.00 100.00
10 91.80 100.00 100.00 100.00
30 96.72 100.00 100.00 100.00
50 97.29 100.00 100.00 100.00
100 96.91 100.00 100.00 100.00
1000 97.03 100.00 100.00 100.00
5000 96.66 100.00 100.00 100.00
10000 96.93 100.00 100.00 100.00
30000 96.83 99.99 100.00 100.00
50000 96.97 100.00 100.00 100.00
Table 15: Percentage IIA Violations Copeland
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 99.95 100.00 100.00 100.00
10 99.99 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 97.30 99.99 100.00 100.00
10 97.75 100.00 100.00 100.00
30 99.80 100.00 100.00 100.00
50 99.99 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 99.99 100.00 100.00 100.00
5000 99.99 100.00 100.00 100.00
10000 99.98 100.00 100.00 100.00
30000 99.97 100.00 100.00 100.00
50000 99.98 100.00 100.00 100.00
Table 16: Percentage IIA Violations Baldwin
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 92.87 99.39 100.00 100.00
10 94.79 99.99 100.00 100.00
30 99.96 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.76 100.00 100.00
30 83.23 99.76 100.00 100.00
50 83.38 99.78 100.00 100.00
100 83.35 99.75 100.00 100.00
1000 82.49 99.76 100.00 100.00
5000 82.63 99.73 100.00 100.00
10000 83.14 99.74 100.00 100.00
30000 82.74 99.74 100.00 100.00
50000 82.87 99.70 100.00 100.00
A.2 Unanimity
Table 17: Percentage U Violations Plurality and Top Two
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 16.23 58.00 86.12 96.61
3 2.56 17.38 41.34 62.30
4 0.43 4.59 13.93 28.63
5 0.05 1.14 4.58 10.87
10 0.00 0.00 0.00 0.07
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 59.94 86.13 96.56
3 11.02 21.64 41.69 63.20
4 4.42 7.96 15.58 28.35
5 2.32 2.83 5.00 10.14
10 0.19 0.06 0.04 0.05
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
Table 18: Percentage U Violations Minimax
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 16.23 42.04 65.70 81.92
3 2.56 8.90 18.46 29.53
4 0.43 1.61 3.72 6.99
5 0.05 0.34 0.73 1.21
10 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 44.53 65.78 82.17
3 11.02 12.93 19.29 30.03
4 4.42 4.04 4.56 6.42
5 2.32 0.98 0.89 1.23
10 0.19 0.01 0.00 0.00
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
A.3 JS Frequencies
Recall that the JS Frequencies for Plurality, Instant Runoff, Minimax, and Top Two are always less than 0.01% in elections with at least three alternatives and two voters. For the rest of the systems, the JS Frequency tables are below.
Table 19: JS Frequencies Borda Count
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 5.16 0.01 0.00 0.00
30 0.11 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.12 0.00 0.00
30 15.48 0.07 0.00 0.00
50 15.71 0.18 0.00 0.00
100 16.15 0.06 0.00 0.00
1000 16.00 0.15 0.00 0.00
5000 15.48 0.14 0.00 0.00
10000 15.37 0.15 0.00 0.00
30000 15.89 0.09 0.00 0.00
50000 15.75 0.14 0.00 0.00
Table 20: JS Frequencies Coombs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 5.16 0.01 0.00 0.00
30 0.02 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.10 0.00 0.00
30 12.81 0.01 0.00 0.00
50 13.96 0.00 0.00 0.00
100 14.50 0.00 0.00 0.00
1000 14.20 0.02 0.00 0.00
5000 13.90 0.02 0.00 0.00
10000 13.40 0.01 0.00 0.00
30000 14.29 0.02 0.00 0.00
50000 14.02 0.01 0.00 0.00
Table 21: JS Frequencies Dowdall
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 0.22 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 7.41 0.02 0.00 0.00
30 3.30 0.00 0.00 0.00
50 3.27 0.00 0.00 0.00
100 3.16 0.00 0.00 0.00
1000 3.13 0.01 0.00 0.00
5000 3.11 0.00 0.00 0.00
10000 2.97 0.00 0.00 0.00
30000 2.98 0.00 0.00 0.00
50000 3.10 0.00 0.00 0.00
Table 22: JS Frequencies Copeland
(a) IC Data
Number of voters Number of alternatives
3       4       5 6
10 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3       4       5 6
10 1.90 0.00 0.00 0.00
30 0.15 0.00 0.00 0.00
50 0.08 0.00 0.00 0.00
100 0.02 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.02 0.00 0.00 0.00
10000 0.03 0.00 0.00 0.00
30000 0.01 0.00 0.00 0.00
50000 0.02 0.00 0.00 0.00
Table 23: JS Frequencies Baldwin
(a) IC Data
Number of voters Number of alternatives
3      4      5 6
10 5.16 0.01 0.00 0.00
30 0.10 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.25 0.01 0.00
30 16.87 0.27 0.00 0.00
50 16.79 0.28 0.00 0.00
100 17.14 0.16 0.00 0.00
1000 17.27 0.22 0.00 0.00
5000 16.63 0.27 0.00 0.00
10000 16.68 0.29 0.00 0.00
30000 17.07 0.22 0.00 0.00
50000 16.88 0.33 0.00 0.00
Table 24: JS Frequencies Pairwise Majority
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 76.19 45.27 21.02 7.50
30 82.69 56.23 31.23 13.97
50 84.64 59.83 35.31 17.34
100 86.92 64.02 39.92 21.08
1000 90.05 70.89 48.37 28.82
5000 90.65 72.55 50.99 30.84
10000 90.81 72.64 51.11 31.40
30000 90.89 73.53 52.06 31.89
50000 91.08 73.77 52.04 32.39
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 84.50 51.12 22.47 7.69
30 90.42 64.15 33.53 14.42
50 91.81 67.86 38.56 17.78
100 92.77 71.49 43.87 21.85
1000 93.54 75.46 52.17 30.32
5000 93.76 75.97 53.24 32.27
10000 93.67 75.87 53.28 32.92
30000 93.55 75.94 53.08 33.02
50000 93.66 76.10 53.51 32.80